August 18, 2021 Calorimeter
'Meeting Day/Time: every other WEDNESDAY at 11:00 a.m. JLab time
- To join via a Web Browser, go to the page  https://bluejeans.com/907185247.
- To join via Polycom room system go to the IP Address: 22.214.171.124 (bjn.vc) and enter the meeting ID: 907185247.
- To join via phone, use one of the following numbers and the Conference ID: 907185247.
- US or Canada: +1 408 740 7256 or
- US or Canada: +1 888 240 2560
- Upon connection all microphones are automatically muted. To unmute your mike on a Polycom or equivalent unit, enter *4. Unmuting on a computer is trivial as there is a microphone button than can be clicked.
- More information on connecting to bluejeans is available.
Participant Direct Lines
- JLab Phone in CC L207 is 757-269-7084 (usual room)
- JLab Phone: in CC F326 is 757-269-6460
- Phone in the Regina Video-conference Suite is 306-585-4204
- Athens Phone: in Christina's office is 011-30-210-727-6947
- FCAL HDFCAL log book
- BCAL HDBCAL log book
- CCAL HDCCAL log book
- Calibrations: 2020 Data Production; RunPeriod-2019-11 Validation; Offline Monitoring Data Validation
- Photon Reconstruction Task Force
Goals for Calorimetry Group
- Determine preliminary photon reconstruction efficiencies as a function of E, phi and theta in data and simulation with a point-to-point precision of at least 5%.
- Measure systematics of pi0/eta mass calibration as a function of detector position to a precision of at least 5 MeV.
- Demonstrate agreement of photon reconstruction efficiency and resolution between data and simulation as a function of E, phi and theta to within 5%.
- Run Preparations
- BCAL LED - LED occupancy runs log entry 
- BCAL (offline)
- FCAL (offline)
- Brief follow up: Churamani
- Action Items
- Any other business
Attendees: Igal, Sean, Jon, Susan, Varun., Malte, Mark, Tolga, Matt, Churamani, Rebecca, Karthik, Zisis
- Start of run has been delayed at least until Sept 4 for Hall D.
- Run Preparations
- FCAL: Malte reported that 80 bases were tested in the EEL building and 25 were re-inserted into the holes of the FCAL. Mark turned them on: 14 ok, 11 have communication issues that will be fixed today. The remaining 55 bases will be spares. Mark mentioned that Nick saw a load based on the firmware's calculations and interrupts which can lead to the lock up; he lightened the 'load' and the new version of the firmware will be tested on 100 bases before uploading it to all bases. The Sep 4 start date provides enough time to do this.
- BCAL: Mark reported on the hot check out: Mark will do chillers, etc, Jon will do the voltages, Varun the SiPM checks.
- CCAL: no report
- Calibration: no report
- BCAL LED - LED occupancy runs log entry shown by Varun. An issue with a module being turned off due to the single-channel pulsing script was fixed by Jon. The occupancy plots look good: 3 dead channels (known from 2020) and potentially a new dead channel. Varun will continue to investigate and then take dark rate runs (he will ask Sergei for DAQ control).
- BCAL (offline): no report
- FCAL (offline): no report
- Reconstruction: Churamani presented an update on the Tag & Probe method for the η'. He introduces post-cuts, numbered 0 to 4, gradually introducing CL, UE, SQ, inv mass and BCAL conditions. He investigated the sharp drop in the yield below the η' mass as a follow up from his last presentation. It turns out that once the η' inv mass cut is used, this effect is not seen. Sl. 14 shows what looks like overlapping distributions, but this vanishes with the inv mass cut. Sean pointed out that it could be some cut inherited from the trees that Churamani is using, but this may not be of interest as the problem does not persists with inv mass and BCAL cuts. A key next step is to test bggen vs data with the same cuts.
- Simulations: no report
- Action Items: the short, long action items and work packages were updated and reviewed. Neutron work will start soon by Tegan et al.
- Any other business